OpenStudy (anonymous):

How do you solve: square root of (x+6)=x

OpenStudy (anonymous):

x=3 and x=-2

OpenStudy (anonymous):

just square both sides, and you will get a quadratic equation, then factor it out.

OpenStudy (gw2011):

square root (x+6) = x Square both sides x+6 = x^2 x^2-x-6 = 0 (x-3)(x+2) = 0 x-3 = 0 x = 3 x+2 = 0 x = -2 Even though x=3 and x=-2 the only solution is x=3 because of the following: square root (3+6) = 3 Square root 9 = 3 which is a true statement square root (-2+6) = square root 4 = 2 but the second solution states that x=-2 and therefore the square root 4 does not equal -2

OpenStudy (anonymous):

-2 is a extraneous root

OpenStudy (anonymous):

it is an extraneous root